If x^1/2 = 1/8, what is the value of x^2/3?
Answer by Alan3354(69443)

If x^1/2 = 1/8, what is the value of x^2/3?
------------
x^(3/6) = 1/8
x^(1/6) = 1/2
==========
x^(2/3) = x^(4/6) = (1/2)^4
x^(2/3) = 1/16